100% ■· Tue Mar 6 12:12 PM Diego Saavedra a is located in Connect Project 13 selis a handmade automobile as its only product. Each automobile is identical, however, they can be distinguished by their unique ID Carlson Auto Dealers Inc. number. At the beginning of 2016, Cartson had three cars in inventory, as follows During 2016, each of the three autos sold for $90,000. Additional purchase. Cited ndronolga, oder) and sales fo·the yer were as Car IDCost Selling Price s 90,000 63,000 64,500 66,000 69,000 0,500 72,000 93,000 215 1 00,500 105,000 2016 ending inventory and cest of goods sold assuming the company uses the specific identification Calculate ending inventary and cost of goods sold assuming FIFO and a periodie invertory system S Calculate ending inventory and cost of goods sold assuming LiFO and a periodic inventory system Calcuiate ending inventory and cost of goods sold assuming the average cost method and a periodic MacBook Air
Answer #1
1 Specific Identification Method
Car ID Cost of goods sold Sales price
203 $60,000 $90,000
207 $60,000 $90,000
210 $63,000 $90,000
211 $63,000 $90,000
212 $63,000 $93,000
214 $66,000 $96,000
215 $69,000 $100,500
217 $72,000 $105,000
218 $72,300 $106,500
TOTAL $588,300 $861,000
ENDING INVENTORY
Car ID Purchase cost
213 $64,500
216 $70,500
219 $75,000
TOTAL $210,000
2016 ending inventory $210,000
2016 cost of goods sold $588,300
2 FIFO and Periodic inventory system
ENDING INVENTORY
Quantity of ending inventory 3
Amount of ending inventory=Cost of last three purchases
Amount of ending inventory=
ID Cost
217 $72,000
218 $72,300
219 $75,000
A Ending Inventory $219,300
B Beginning inventory $        183,000 (60000+60000+630000)
Purchased during the year
Cae ID Cost
211 $63,000
212 $63,000
213 $64,500
214 $66,000
215 $69,000
216 $70,500
217 $72,000
218 $72,300
219 $75,000
C Total purchase during the year $615,300
D=B+C Inventory available $        798,300
E=D-A Cost ofgoods sold $        579,000
2016 ending inventory $219,300
2016 cost of goods sold $        579,000
3 LIFO and Periodic inventory system
ENDING INVENTORY
Quantity of ending inventory 3
Amount of ending inventory=Cost of first three cars
Amount of ending inventory:
Cost
$60,000
$60,000
$63,000
Total $183,000
A Ending Inventory $183,000
B Beginning inventory $        183,000 (60000+60000+630000)
Purchased during the year
Cae ID Cost
211 $63,000
212 $63,000
213 $64,500
214 $66,000
215 $69,000
216 $70,500
217 $72,000
218 $72,300
219 $75,000
C Total purchase during the year $615,300
D=B+C Inventory available $        798,300
E=D-A Cost ofgoods sold $        615,300
2016 ending inventory $183,000
2016 cost of goods sold $        615,300